A 2-kg ball B traveling at 22m.s-1 Overtakes a 4-kg ball traveling in the same direction at 10m.s-1 Find the velocity of each ball after collision.


1
Expert's answer
2021-04-20T17:30:15-0400

"\begin{cases}
 \frac{m_bv_b^2}{2}+\frac{m_av_a^2}{2}=\frac{m_bu_b^2}{2}+\frac{m_au _a^2}{2},\\
 m_bv_b+m_av_a=m_bu_b+m_au_a;
\end{cases}"

"\begin{cases}
 m_b(v_b^2-u_b^2)=m_a(u_a^2-v_a^2), \\
 m_b(v_b-u_b)=m_a(u_a-v_a);
\end{cases}"

"\begin{cases}
 v_b+u_b=u_a+v_a, \\
 v_b-v_a=u_a-u_b;
\end{cases}"

"\begin{cases}
 u_a=\frac{2m_bv_b+(m_a-m_b)v_a}{m_a+m_b}, \\
 u_b=\frac{2m_av_a+(m_b-m_a)v_b}{m_a+m_b};
\end{cases}"

"\begin{cases}
 u_a=18~\frac ms, \\
 u_b=6~\frac ms.
\end{cases}"